OpenStudy (anonymous):

The highway mileage m in miles per gallon for a compact car is approximated by m(s) = -0.025s2 + 2.45s - 30, where s is the speed in miles per hour. What is the maximum mileage for this compact car to the nearest tenth of a mile per gallon? __________ mi/gal. (Enter the number only.) To the nearest whole number, what speed results in this mileage? __________ mi/h (Enter the number only.)

OpenStudy (anonymous):

differentiate with respect to s and equate it to 0. you will get the speed for which the mileage is maximum. use that to calculate the mileage.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do i do it to all the s's or the end one?

OpenStudy (anonymous):

huh? do you know how to differentiate? is this a calculus question?

OpenStudy (anonymous):

its a lagebra question

OpenStudy (anonymous):

have you done calculus before?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i see. this is going to be hard without using derivatives. have you done parabolas in geometry?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

well? this is a down facing parabola. its vertex is at -b/2a that is the speed which gives maximum mileage.

OpenStudy (anonymous):

so speed = -b/2a = -2.45/(2 x - 0.025) = -2.45/-0.05 = 2.45/0.05 = 245/5 = 49 mph

OpenStudy (anonymous):

substitute that in the mileage equation and get the maximum mileage.

OpenStudy (anonymous):

ok